# Dilaton Effective Field Theory across the Conformal Edge

### v1

@@ Acknowledgements.
We thank Ed Bennett and Biagio Lucini for useful discussions. The work of M.P. has been supported in part by the STFC Consolidated Grants No. ST/T000813/1 and ST/X000648/1. M.P. received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program under Grant Agreement No. 813942. J.I. has been supported by STFC under grants No. ST/X000745/1 and No. ST/X003167/1. Research Data Access Statement—The data generated for this manuscript can be downloaded from Ref. [data_release]. Open Access Statement—For the purpose of open access, the authors have applied a Creative Commons Attribution (CC BY) license to any Author Accepted Manuscript version arising.
Table 3: Numerical lattice measurements taken from Table IX of Ref. [LatticeStrongDynamics:2023bqp] with $\beta=4.8$ (see also the data release [SU(3)-data_release]), for the $SU(3)$ theory coupled to $N_{f}=8$ fermions transforming in the fundamental representation. The errors in parenthesis includes statistical and systematics, but correlations are neglected. The fermion mass, amam, appears as a parameter in the lattice staggered action, and hence there is no associated uncertainty.
@@
Table 4: Numerical lattice measurements for the $SU(2)$ theory coupled to $N_{f}=1$ fermion transforming in the adjoint representation [Athenodorou:2021wom, SU(2)-1-data_release, SU(2)-2-data_release], for ensembles with $\beta=2.1$, and the largest available volumes. The errors in parenthesis are statistical only, and correlations are neglected. These numerical results are obtained with Wilson fermions, hence a measure of symmetry breaking effects due to finite fermion mass is provided by the PCAC mass, $am_{\rm PCAC}$, extracted according to the process described in Ref. [DelDebbio:2007wk].
@@ Appendix A Maximum likelihood analysis and lattice data
We here collect some technical details needed to reproduce our main results. First, we summarize in Tables 3 and 4 the spectroscopic measurements for pNGBs and dilaton, taken from Refs. [LatticeStrongDynamics:2023bqp] and [Athenodorou:2021wom]. These measurements are combined with Eqs. (9), (10), and (11), to define the $\chi^{2}$, in the following way.

### v2

@@ Acknowledgements.
We thank Ed Bennett and Biagio Lucini for useful discussions. The work of M.P. has been supported in part by the STFC Consolidated Grants No. ST/T000813/1 and ST/X000648/1. M.P. received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program under Grant Agreement No. 813942. J.I. has been supported by STFC under grants No. ST/X000745/1 and No. ST/X003167/1. Research Data Access Statement—The data generated for this manuscript can be downloaded from Ref. [15]. Open Access Statement—For the purpose of open access, the authors have applied a Creative Commons Attribution (CC BY) license to any Author Accepted Manuscript version arising.
Table 3: Numerical lattice measurements taken from Table IX of Ref. [26] with $\beta=4.8$ (see also the data release [39]), for the $SU(3)$ theory coupled to $N_{f}=8$ fermions transforming in the fundamental representation. The errors in parenthesis includes statistical and systematics, but correlations are neglected. The fermion mass, amam, appears as a parameter in the lattice staggered action, and hence there is no associated uncertainty.
@@
Table 4: Numerical lattice measurements for the $SU(2)$ theory coupled to $N_{f}=1$ fermion transforming in the adjoint representation [22, 18, 17], for ensembles with $\beta=2.1$, and the largest available volumes. The errors in parenthesis are statistical only, and correlations are neglected. These numerical results are obtained with Wilson fermions, hence a measure of symmetry breaking effects due to finite fermion mass is provided by the PCAC mass, $am_{\rm PCAC}$, extracted according to the process described in Ref. [35].
@@ Appendix A Maximum likelihood analysis and lattice data
We here collect some technical details needed to reproduce our main results. First, we summarize in Tables 3 and 4 the spectroscopic measurements for pNGBs and dilaton, taken from Refs. [26] and [22]. These measurements are combined with Eqs. (9), (10), and (11), to define the $\chi^{2}$, in the following way.